Circular No. 8295

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ac IN IC 4769 L. A. G. Monard, Pretoria, S. Africa, reports the discovery of an ap-

parent supernova (mag ∼ 16.9 ± 0.3) on unfiltered CCD images taken on Feb. 29.104 UT and confirmed on Mar. 1.115 (mag 16.8). SN 2004ac is located at α = 18^{h}47^{m}43^{s}.67, δ = −63^{o}08′54′′.6 (equinox 2000.0), which is 9′′ west and 31′′ north of the nucleus of IC 4769. The new object is not visible on a red image from the Digitized Sky Survey (limiting mag 19.5) or on earlier images taken by Monard on 2003 Sept. 18.77 (limiting mag 18.7) and Oct. 16.75 (limiting mag 18.5).

2004 DW D. Rabinowitz and S. Tourtellotte, Yale University — in collaboration

with M. Brown, California Institute of Technology; and C. Trujillo, Gemini Observatory — report the following magnitudes for 2004 DW (cf. IAUC 8291) measured with the SMARTS 1.3-m telescope at Cerro Tololo on Feb. 26: B = 19.70 ± 0.01, V = 19.01 ± 0.03, R = 18.64 ± 0.02, I = 18.25 ± 0.04. As the values for B–V, V–R, and V–I are close to solar values, the spectral reflectance is neutral.

IRAS 05436–0007 G. Maheswar and H. C. Bhatt, Indian Institute of Astrophysics, Ban-

galore, report on observations of the optical nebula associated with IRAS 05436–0007 (cf. IAUC 8284) made at the Kavalur Observatory 2.3-m Vainu Bappu Telescope (VBT) and the Hanle Observatory 2-m Himalayan Chandra Telescope (HCT). VBT R-band CCD imaging on Feb. 19.7 UT yields magnitudes 16.8 in a 6′′ aperture and 13.6 in an 80′′ aperture. HCT CCD imaging on Feb. 22.7 gave V = 18.1, R = 16.5, and I = 14.5 in a 6′′ aperture, and V = 14.0, R = 13.2, and I = 12.1 in an 80′′ aperture. An optical spectrum (resolution of 0.12 nm/pixel) with the HCT (+ HFOSC) on Feb. 23.7 shows Hα and near-infrared Ca II triplet lines in emission. Hα shows a P-Cyg profile with an eqivalent width of the emission component of 3.2 nm, the absorption component having a peak at a velocity of 460 km/s. The O I line at 777.3 nm is seen in absorption with an equivalent width of 0.24 nm.

2004 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green